Find a polynomial for the sum of the areas of a circles if the circle has a fadius of 17 and the other has a radius of 8
Answer by checkley77(12844)

AREA=PIR^2
AREA=3.14*17^2+3.14*8^2
AREA=3.14*289+3.14*64
AREA=3.14(289+64
AREA=3.14*353
AREA=1,108.42 ANS.